APACHE POINT OBSERVATORY SDSS 2.5M OBSERVING LOG Sunday August 8, 2010 (MJD 55417) ---=== OBSERVING TEAM ===--- Dmitry Bizyaev Dan Oravetz Paul Harding Nathan DeLee ---=== OBSERVING PLAN ===--- Post-shutdown Tasks/Tests ---=== OBSERVING SUMMARY ===--- The night was lost to clouds and humidity. We reset the nominal MIG values in the mirror numbers to provide a new baseline following the home/relax procedure from the other night. We tested accessing MJCI and MJDD with the aliased 'marvels' in a Terminal window and it appears there are connection issues. Nathan was able to get things working with a modified host name (marvels-p as opposed to marvels), but he will talk to Fritz and Jon about resolving the problem so things are back to normal from a procedural point of view. MARVELS personnel began work on repairing the broken fibers in the fiber link. Elena M. wrote a script to complete the grid of engineering camera observerations called ecamOffsetsTest. Paul worked on gcameraICC software. ---=== OBSERVING LOG ===--- Afternoon ========= Arrived to cloudy skies with threats of precipitation. We tested opening MJCI and MJDD with Nathan's oversight and there are problems getting things to talk properly. Nathan used a different host name (e.g. marvels-p) and it worked, but he will talk to Fritz and Jon to resolve the problem because as it stands we cannot use the aliased 'marvels' in the Terminal to get the proper connections in the MJCI window that pops up. MARVELS folks were beginning work on repairing the broken fibers along the fiber link. The testActors script only produced the following output: alerts ping ** FAILED ** alerts status ** FAILED ** apo ping ** FAILED ** apo status ** FAILED ** With actors offline the getVersions script didn't provide many versions. Disk space: [observer@hub25m-p ~]$ df -h Filesystem Size Used Avail Use% Mounted on /dev/sda2 95G 15G 75G 17% / /dev/sda4 793G 92G 661G 13% /home /dev/sda1 99M 19M 75M 20% /boot tmpfs 5.9G 0 5.9G 0% /dev/shm /dev/drbd1 13T 2.1T 11T 17% /data Night ===== Checked out the telescope. The telescope moved in all three axes normally. The fiducials were still only reading in altitude, but this isn't news and work continues on getting all fiducials to read properly. The mirror numbers are still not within spec as a result of the homing/relaxing procedure completed the other night. 03:00Z - Elena wrote a script to complete the engineering camera grid of exposures. 04:10Z - Reset the nominal MIG values in the mirror numbers. 07:49Z - Paul worked on gcameraICC software and got the overscan to work. 07:54Z - Called it a night with the clouds/approaching clouds and humidity still at work. 07:55Z - endNight rejoiced. ---=== IMAGING RUN SUMMARY ===--- Run Time Stripe Lambda Last Flavor Comments Start End Begin End Frame ------------------------------------------------------------------------- ---=== IMAGING RUN DETAILS ===--- ---=== SKIPPY RESULTS ===--- Run Frame nFrames stars muErr muRms nuErr nuRms rot az el --------------------------------------------------------------------------- ---=== LTMATCH RESULTS ===--- Run Field nFields alt az nGood rowMean rowSig colMean colSig rot ------------------------------------------------------------------------ ---=== SPECTROSCOPY DATA SUMMARY ===--- Summary Checked (y/n): n QA Procedures Done (y/n): n UT Exp Time flavor comment (S/N)^2 totals ========================================== b1 r1 b2 r2 ---=== TELESCOPE OFFSETS AND SCALE I ===--- Time Instrument Az Alt Rot Scale pos offset pos offset pos offset ------------------------------------------------------------------------------ ---=== TELESCOPE OFFSETS AND SCALE II ===--- ---=== FOCUS LOG ===--- setmir piston Temp Wind Time Inst scale M1 M2 Foc Az Alt (C) MPH Dir filt fwhm ------------------------------------------------------------------------------ ---=== WEATHER LOG ===--- Wind Time Temp F Dewp F MPH Direction Dust DIMM Sky 00:24Z 64 47 4 276 (W) 371 - clouds 00:57Z 63 48 1 308 (NW) 367 - 01:30Z 60 50 8 10 (N) 1170 - 02:03Z 58 51 11 66 (ENE) 520 - 02:36Z 58 50 10 64 (ENE) 477 - 03:09Z 58 49 5 92 (E) 403 - 03:42Z 58 49 6 83 (E) 332 - 04:15Z 52 51 19 133 (SE) 622 - 04:48Z 52 52 16 115 (ESE) 535 - 05:21Z 51 51 20 129 (SE) 764 - 05:54Z 52 50 21 134 (SE) 766 - 06:26Z 53 48 14 150 (SSE) 334 - 06:59Z 53 48 8 80 (E) 561 - 07:32Z 54 47 2 124 (SE) 253 - ---=== TELESCOPE STATUS ===--- The engineering camera is mounted with the telescope at stow. The LN2 autofill lines are not connected. ---=== SOFTWARE USED ===--- This information automatically included upon opening logTool: IOP/SOP: v4_40_0 Watcher: v2_36_0 MCP: v6_6_5 TPM: tpm_v3_5_0 AstroDa: v15_13_1 TCC: 1067750 packets sent successfully sdssProcedures: SoS: plate-mapper: This information taken from getVersions script: STUI: 1.0b2 2010-07-21 alerts: ** FAILED ** apo: ** FAILED ** boss: trunk+svn115445M boss.daq: 1.2.14 boss.specMech: Offline gcamera: trunk+svn115470 guider: branch_dstn-fiber-centers+svn115457M hub: Tag: UNTAGGED_REVISION 502:504M mcp: v6_6_5 mcp.plc: 25 platedb: ** FAILED ** sop: ** FAILED ** sos: ** FAILED ** ---=== MIRROR NUMBERS ===--- Mirror numbers from the checkout before resetting NOMINAL MIG values: PRIMARY: -------- Scale: 1.000000 MIGS TONIGHT NOMINAL Axial A -5.4580 -0.6940 Axial B 16.0490 -0.0410 Axial C 15.9110 -0.3710 Trans D 11.5860 11.6170 Lateral E 10.3505 10.3505 Lateral F 11.9253 11.9253 GALILS Commanded: 9900. 400. 5850. 1200. 2950. 2750. Actual: 9917. 401. 5832. 1205. 2954. 2757. SETMIR VALUES PriDesOrient: 0.00 -3.30 5.10 278.08 178.06 PriOrient: 0.00 -3.04 5.26 277.70 178.07 SECONDARY: ---------- Focus: 0.00 Air Temp.: 14.7 Alt.: 29.999993 MIGS TONIGHT NOMINAL Axial A -3.8330 18.6150 Axial B 19.5500 15.6120 Axial C 17.8190 13.8170 Trans D 14.1290 -11.7780 GALILS Commanded: -3453999. -3472733. -3440437. -1800. 100. Actual: -3454009. -3472664. -3440479. -1778. 95. SETMIR VALUES SecDesOrient: -2743.00 -9.00 -4.00 -58.00 23.67 SecOrient: -2743.00 -9.03 -4.02 -58.30 23.86 Mirror numbers having reset the NOMINAL MIG values: PRIMARY: -------- Scale: 1.000000 MIGS TONIGHT NOMINAL Axial A -5.4580 -5.4580 Axial B 16.0490 16.0490 Axial C 15.9120 15.9120 Trans D 11.5860 11.5860 Lateral E 10.3505 10.3505 Lateral F 11.9253 11.9253 GALILS Commanded: 9900. 400. 5850. 1200. 2950. 2750. Actual: 9917. 401. 5832. 1205. 2954. 2757. SETMIR VALUES PriDesOrient: 0.00 -3.30 5.10 278.08 178.06 PriOrient: 0.00 -3.04 5.26 277.70 178.07 SECONDARY: ---------- Focus: 0.00 Air Temp.: 11.7 Alt.: 30.001743 MIGS TONIGHT NOMINAL Axial A -3.8330 -3.8330 Axial B 19.5520 19.5510 Axial C 17.8200 17.8200 Trans D 14.1300 14.1300 GALILS Commanded: -3454025. -3472639. -3440500. -1800. 100. Actual: -3454009. -3472664. -3440479. -1778. 95. SETMIR VALUES SecDesOrient: -2743.00 -9.00 -4.00 -58.00 23.67 SecOrient: -2743.00 -8.99 -3.99 -58.28 23.84 ---=== PROBLEMS IN DETAIL ===---